livingswag413:

Find the length of the base of a square pyramid if the volume is 576 cubic inches and has a height of 3 inches. 4 inches 8 inches 16 inches 24 inches This is all i have Thank You So Much

justus:

Volume of a square pyramid = area of base *height / 3 576 = l^2 *3/3 576 = l^2 l = sqrt576 l = 24 inches l is length

livingswag413:

THANK YOU SO MUCH!!!

justus:

No problem.

justus:

Good luck

livingswag413:

Wishing The Same Back Your Way!!

justus:

:)

justus:

If you need more help, Im here

livingswag413:

:'))

livingswag413:

You're literally a lifesaver.

justus:

Thank you, I appreciate that.

livingswag413:

i wouldve been kicked out loll and math has always been a problem for me

livingswag413:

Honestly though

justus:

Ah I see, I don't like math either.

justus:

Im not the best at math

livingswag413:

oh

justus:

But I hope you pass the class!

livingswag413:

Thank You!!

justus:

Yup!